OpenStudy (anonymous):

How many solutions does the equation 2|x − 3| + 4 = 4 have? Zero One Two Infinitely Many

OpenStudy (hoblos):

2|x − 3| + 4 = 4 2|x-3| = 0 |x-3| = 0 thus x=3 is the only solution

OpenStudy (anonymous):

so it's zero

OpenStudy (hoblos):

you have one solution... so it's One

OpenStudy (anonymous):

oh ok